OpenStudy (anonymous):

(42b3+9b^2+42b+33)/(6b+3)

jhonyy9 (jhonyy9):

(42b(b2 +1) +3(3b2 +11))/3(2b+1) =3(14b(b2 +1) +(3b2 +11))/3(2b+1) = =(14(b2 +1) +3b2 +11)/(2b+1)